OpenStudy (anonymous):

what is the area of a circle with a circumference of 43.98226 inches 3.14159 for π

OpenStudy (jdoe0001):

\(\bf \underline{circumference}\textit{ of a circle}=2\pi r\qquad \begin{cases} circumference = &43.98226\\ \pi =&3.14159 \end{cases} \\ \quad \\ circumference =2\pi r\implies \cfrac{circumference}{2\pi }=r \\ \quad \\ \cfrac{43.98226}{2\cdot 3.14159}={\color{brown}{ r}} \\ \quad \\ \quad \\ \underline{area}\textit{ of a circle}=\pi {\color{brown}{ r}}^2\)

OpenStudy (jdoe0001):

once you get "r" or the radius use that in the "area" equation